# On the complexity of matrix Putinar’s Positivstellensatz

### v1

@@ 2.1. Notations
2.2. Some basics in real algebraic geometry
@@
(2.1) $\mbox{QM}[G]_{k}^{\ell}=\left\{S+\sum_{i=1}^{t}P_{i}^{T}GP_{i}\mid t\in\mathbb{N},~{}S\in\Sigma[x]^{\ell}_{k},~{}P_{i}\in\mathbb{R}[x]_{\lceil k-\deg(G)\rceil/2}^{m\times\ell}\right\}.$
Note that $\Sigma[x]^{\ell}\subseteq\mbox{QM}[G]^{\ell}$. For the case $\ell=1$, we simply denote that
$\Sigma[x]_{k}:=\Sigma[x]_{k}^{1},~{}\mbox{QM}[G]_{k}:=\mbox{QM}[G]_{k}^{1}.$
The quadratic module $\mathrm{QM}[G]^{\ell}$ is said to be archimedean if there exists $r>0$ such that
$\left(r-\|x\|_{2}^{2}\right)I_{\ell}\in\mathrm{QM}[G]^{\ell}.$
The polynomial matrix $G$ determines the semialgebraic set
@@
If $\mbox{QM}[G]^{\ell}$ is archimedean, then the set $K$ must be compact. However, when $K$ is compact, it is not necessarily that $\mbox{QM}[G]^{\ell}$ is archimedean. The following is matrix Putinar’s Positivstellensatz.
@@
Suppose that $G$ is an $m$-by-$m$ symmetric polynomial matrix such that $\mbox{QM}[G]^{\ell}$ is archimedean. If the $\ell$$-\ell$ symmetric polynomial matrix $F\succ 0$ on $K$, then $F\in\mbox{QM}[G]^{\ell}$.
2.3. Some technical propositions
In this subsection, we present two useful propositions that are used in our proofs. The following is the classical Lojasiewicz inequality.
@@ 2.3. Some technical propositions
Let $A$ be a closed and bounded semialgebraic set and $f$ and $g$ two continuous semialgebraic functions from $A$ to $\mathbb{R}$ such that $f^{-1}(0)\subset g^{-1}(0)$. Then there exist two positive constants $\eta$, $\kappa$ such that:
@@
The smallest constant $\eta$ is called the Lojasiewicz exponent and $\kappa$ is the corresponding Lojasiewicz constant.
@@
3. Degree bounds for scalar polynomial inequalities
In this section, we consider the case that $G(x)$ is a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$. Then, the set $K$ as in (1.1) can be equivalently described as
(3.1) $K=\{x\in\mathbb{R}^{n}\mid g_{1}(x),\dots,\dots,g_{m}(x)\geq 0\}.$
If the symmetric polynomial matrix $F(x)\succ 0$ on $K$, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (1.7). It improves the exponential bound shown by Helton and Nie [14]. The proof generalizes the techniques recently developed in [2, 3] to the matrix case.
@@ 3. Degree bounds for scalar polynomial inequalities
$\bar{\Delta}:=\left\{x\in\mathbb{R}^{n}\mid\sqrt{n}-x_{1}-\cdots-x_{n}\geq 0,1+x_{1}\geq 0,\ldots,1+x_{n}\geq 0\right\}.$
@@
$B_{\alpha}^{(t)}(x)=\frac{t!}{\alpha_{1}!\cdots\alpha_{n}!}(n+\sqrt{n})^{-t}\left(\sqrt{n}-x_{1}-\cdots-x_{n}\right)^{t-|\alpha|}\left(1+x_{1}\right)^{\alpha_{1}}\cdots\left(1+x_{n}\right)^{\alpha_{n}},$
@@
$\|F\|_{B,t}=\max\limits_{\alpha\in\mathbb{N}^{n}_{t}}\{\|F_{\alpha}^{(t)}\|_{2}\}.$
For convenience of notations, we denote $\|F\|_{B,d}$ by $\|F\|_{B}$. Similarly, $F$ can also be expressed in the monomial basis $\{x^{\alpha}:\alpha\in\mathbb{N}^{n}_{d}\}$ as
@@
The spectral norm of $F$ is defined as
$\|F\|_{2}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d}}\{\|F_{\alpha}\|_{2}\}.$
Suppose that $p\in\mathbb{R}[x]$ is a homogeneous polynomial with $\deg(p)=d$. The Polya’s Positivstellensatz [36, 37] states that if $p\geq p_{\min}>0$ on the standard simplex $\Delta^{n}$, then the polynomial $(x_{1}+\cdots+x_{n})^{k}p(x)$ has only positive coefficients if
$k>d(d-1)\frac{\|p\|}{2p_{\min}}-d.$
This result has been generalized to the matrix case by Scherer and Hol [43]. It was shown in [43, Theorem 3] that if $F(x)$ is an $\ell$$-\ell$ homogeneous symmetric polynomial matrix with $\deg(F)=d$ and $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $\triangle^{n}$, then for
$k>d(d-1)\frac{\|F\|_{2}}{2F_{\min}}-d,$
all coefficient matrices of $(x_{1}+\cdots+x_{n})^{k}F(x)$ in the monomial basis are positive definite.
In the following, we present a dehomogenized version of matrix Polya’s Positivstellensatz in terms of the Bernstein basis.
@@
Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $\bar{\Delta}$, then for
@@
$\begin{array}[]{ll}\sqrt{n}-\sum\limits_{i=1}^{n}x_{i}&=\sqrt{n}-\sum\limits_{i=1}^{n}((n+\sqrt{n})y_{i}-1)\\ &=\sqrt{n}+n-(n+\sqrt{n})\sum\limits_{i=1}^{n}y_{i}\\ &=(\sqrt{n}+n)(1-\sum\limits_{i=1}^{n}y_{i})\\ &\geq 0,\\ \end{array}$
which implies that $x:=(x_{1},\dots,x_{n})\in\bar{\Delta}$. By computations, we know that for $(y_{1},\dots,y_{n+1})\in\Delta^{n+1}$, the following holds
$\widehat{F}(y_{1},\dots,y_{n+1})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}^{(d)}B_{\alpha}^{(d)}(x)=F(x)\succeq F_{\min}\cdot I_{m}.$
Hence, $\widehat{F}\succeq F_{\min}I_{m}$ on $\Delta^{n+1}$. It follows from Theorem 3 [43] that for $k>d(d-1)\frac{\|\widehat{F}\|_{2}}{2F_{\min}}-d$, all coefficient matrices of $(y_{1}+\cdots+y_{n+1})^{k}\widehat{F}(y_{1},\dots,y_{n+1})$ in the monomial basis are positive definite, i.e., there exist $P_{\alpha}\succ 0$ such that
@@
into the above identity, we get that
@@
Since $\|\widehat{F}\|_{2}=\|F\|_{B}$, the conclusion follows. ∎∎
@@
Then, we have that
(3.2) $\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}.$
Furthermore, there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that
(3.3) $\|P^{(\xi_{0})}\|_{B,d_{P}}=\|P\|_{B}.$
Proof.
Suppoose that $P$ can be expressed as
@@
$\begin{array}[]{ll}\|P^{(\xi)}\|_{B,d_{P}}=\|\xi^{T}P(x)\xi\|_{B,d_{P}}&=\|\sum\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\xi^{T}P_{\alpha}^{(d_{P})}\xi B_{\alpha}^{(d_{P})}(x)\|_{B,d_{P}}\\ &=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi^{T}P_{\alpha}^{(d_{P})}\xi|\}\\ &\leq\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{\|P_{\alpha}^{(d_{P})}\|_{2}\}=\|P\|_{B}.\end{array}$
Thus, we have
$\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}.$
Since there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that
$\|P\|_{B}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|\},$
@@
$\|P^{(\xi_{0})}\|_{B,d_{P}}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|\}=\|P\|_{B}.$
@@
(3.4) $\mathcal{G}(x)=-\min\left(\frac{g_{1}(x)}{\left\|g_{1}(x)\right\|_{B}},\ldots,\frac{g_{m}(x)}{\left\|g_{m}(x)\right\|_{B}},0\right).$
@@
If $\mathcal{G}(x)=0$, then we have $x\in K$ and $d(x,K)=0$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Lojasiewicz inequality holds:
@@
In the following, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (1.7) when $K$ is given by finitely many scalar polynomial inequalities, which improves the exponential bound as in [14].
@@
Let $G(x)$ be a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$ and let $K$ be as in (3.1). Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$. If $F(x)\succeq F_{\min}\cdot I_{\ell}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
@@
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (3.5) and $C>0$ is a constant independent of $F$, $g_{1},\dots,g_{m}$.
Proof.
Without loss of generality, we assume that $\left\|g_{i}\right\|_{B}=1$ $(i=1,\dots,m)$. Otherwise, we can scale $g_{i}$ by $1/\|g_{i}\|_{B}$ and the bound will not change. Let
@@
For the above $\delta$, $v$, it follows from Proposition 3.2 [3] that there exist $h_{i}\in\mathbb{R}[x]$ such that for $x\in\bar{\Delta}$, we have
@@
In the following, we show that $P\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$. Equivalently, for each $\xi\in\mathbb{R}^{\ell}$ with $\|\xi\|_{2}=1$, we prove that for every $x\in\bar{\Delta}$, the following holds

### v2

@@ 2.1 Notations
2.2 Some basics in real algebraic geometry
@@
(2.1) $\mbox{QM}[G]_{k}^{\ell}=\left\{S+\sum_{i=1}^{t}P_{i}^{T}GP_{i}\mid t\in\mathbb{N},~{}S\in\Sigma[x]^{\ell}_{k},~{}P_{i}\in\mathbb{R}[x]_{\lceil k/2-\deg(G)/2\rceil}^{m\times\ell}\right\}.$
Note that $\Sigma[x]^{\ell}\subseteq\mbox{QM}[G]^{\ell}$. For the case $\ell=1$, we simply denote
$\Sigma[x]_{k}:=\Sigma[x]_{k}^{1},~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}~{}\mbox{QM}[G]:=\mbox{QM}[G]^{1}},~{}\mbox{QM}[G]_{k}:=\mbox{QM}[G]_{k}^{1}.$
The quadratic module $\mathrm{QM}[G]^{\ell}$ is said to be Archimedean if there exists $r>0$ such that $\left(r-\|x\|_{2}^{2}\right)\cdot I_{\ell}\in\mathrm{QM}[G]^{\ell}.$ The polynomial matrix $G$ determines the semialgebraic set
@@
If $\mbox{QM}[G]^{\ell}$ is Archimedean, then the set $K$ must be compact. However, when $K$ is compact, it is not necessarily that $\mbox{QM}[G]^{\ell}$ is Archimedean (see [26]). The following is known as the matrix-version of Putinar’s Positivstellensätz.
@@
Suppose that $G$ is an $m$-by-$m$ symmetric polynomial matrix and $\mbox{QM}[G]^{\ell}$ is Archimedean. If the $\ell$$-\ell$ symmetric polynomial matrix $F\succ 0$ on $K$, then we have $F\in\mbox{QM}[G]^{\ell}$.
2.3 Some technical propositions
In this subsection, we present several useful propositions that are used in our proofs. The following is the classical Łojasiewicz inequality.
@@ 2.3 Some technical propositions
Let $A$ be a closed and bounded semialgebraic set, and let $f$ and $g$ two continuous semialgebraic functions from $A$ to $\mathbb{R}$ such that $f^{-1}(0)\subseteq g^{-1}(0)$. Then there exist two positive constants $\eta$, $\kappa$ such that:
@@
The smallest constant $\eta$ is called the Łojasiewicz exponent and $\kappa$ is the corresponding Łojasiewicz constant.
@@
The following theorem is a slight variation of [14, Theorem 3.3], which gives explicit Łojasiewicz exponents for basic closed semialgebraic sets.
Theorem 2.4 ([14]).
Suppose that $g_{1},\dots,g_{s},h_{1},\dots,h_{l}\in\mathbb{R}[x]_{d}$ and let
$S:=\{x\in\mathbb{R}^{n}\mid h_{1}(x)=\cdots=h_{l}(x)=0,g_{1}(x)\geq 0,\ldots,g_{s}(x)\geq 0\}\neq\emptyset.$
Then for any compact set $K\subset\mathbb{R}^{n}$, there is a constant $c>0$ such that for all $x\in K$,
$d(x,S)^{\mathcal{R}(n+l+s-1,d+1)}\leq-c\cdot\min\{g_{1}(x),\dots,g_{s}(x),h_{1}(x),-h_{1}(x),\dots,h_{l}(x),-h_{l}(x),0\},$
where $d(x,S)$ is the Euclidean distance to the set $S$ and $\mathcal{R}(n,d)=d\cdot(3d-3)^{n-1}$.
3 Degree bounds for scalar polynomial inequalities
In this section, we consider the case where $G(x)$ is a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$. Then, the set $K$ as in (1.1) can be equivalently described as
(3.1) $K=\{x\in\mathbb{R}^{n}\mid g_{1}(x)\geq 0,\dots,\dots,g_{m}(x)\geq 0\}.$
If the symmetric polynomial matrix $F(x)\succ 0$ on $K$, we prove a polynomial bound on the degrees of terms appearing in matrix Putinar’s Positivstellensätz. This improves the exponential bound shown by Helton and Nie [15]. The proof essentially generalizes the idea and strategies used in [2, 3] to the matrix case. We refer to Step 1 in the section ‘Outlines of proofs’ for an overview of the proof.
@@ 3 Degree bounds for scalar polynomial inequalities
$\bar{\Delta}=\left\{x\in\mathbb{R}^{n}\mid\sqrt{n}-x_{1}-\cdots-x_{n}\geq 0,1+x_{1}\geq 0,\ldots,1+x_{n}\geq 0\right\}.$
@@
(3.2) $B_{\alpha}^{(t)}(x)=\frac{t!}{\alpha_{1}!\cdots\alpha_{n}!}(n+\sqrt{n})^{-t}\left(\sqrt{n}-x_{1}-\cdots-x_{n}\right)^{t-|\alpha|}\left(1+x_{1}\right)^{\alpha_{1}}\cdots\left(1+x_{n}\right)^{\alpha_{n}},$
@@
$\|F\|_{B,t}=\max\{\|F_{\alpha}^{(t)}\|_{2}:~{}\alpha\in\mathbb{N}^{n}_{t}\}.$
For convenience, we denote $\|F\|_{B,d}$ simply as $\|F\|_{B}$. Similarly, $F$ can also be expressed in the monomial basis $\{x^{\alpha}:\alpha\in\mathbb{N}^{n}_{d}\}$ as
@@
The spectral norm of $F$ is defined as $\|F\|_{2}=\max\{\|F_{\alpha}\|_{2}:~{}\alpha\in\mathbb{N}^{n}_{d}\}.$ For $f(x)\in\mathbb{R}[x]_{d_{1}}$, $g(x)\in\mathbb{R}[x]_{d_{2}}$ and $\ell_{1}\geq\ell_{2}\geq d_{1}$, the following is a basic property of the Bernstein norm (see [13], [3, Lemma 1.5]):
(3.3) $\max\limits_{s\in\bar{\Delta}}|f(s)|\leq\|f\|_{B,\ell_{1}}\leq\|f\|_{B,\ell_{2}},~{}~{}\|fg\|_{B,d_{1}+d_{2}}\leq\|f\|_{B,d_{1}}\|g\|_{B,d_{2}}.$
Suppose that $p(x)$ is a homogeneous polynomial with $\deg(p)=d$. Polya’s Positivstellensätz [40, 41] states that if $p(x)\geq p_{\min}>0$ on the standard simplex $\Delta^{n}$, then the polynomial $(x_{1}+\cdots+x_{n})^{k}p(x)$ has only positive coefficients if $k>d(d-1)\frac{\|p\|}{2p_{\min}}-d.$ This result has been generalized to the matrix case by Scherer and Hol [48]. It was shown in [48, Theorem 3] that if $F(x)$ is an $\ell$$-\ell$ homogeneous symmetric polynomial matrix with $\deg(F)=d$ and $F(x)\succeq F_{\min}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{\ell}}\succ 0$ on $\triangle^{n}$, then for $k>d(d-1)\frac{\|F\|_{2}}{2F_{\min}}-d,$ all coefficient matrices of $(x_{1}+\cdots+x_{n})^{k}F(x)$ in the monomial basis are positive definite.
In the following, we present a dehomogenized version of matrix Polya’s Positivstellensätz in terms of the Bernstein basis.
@@
Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F(x)\succeq F_{\min}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{\ell}}\succ 0$ on $\bar{\Delta}$, then for
@@
$\begin{array}[]{ll}\sqrt{n}-\sum\limits_{i=1}^{n}x_{i}=\sqrt{n}-\sum\limits_{i=1}^{n}((n+\sqrt{n})y_{i}-1)&=\sqrt{n}+n-(n+\sqrt{n})\sum\limits_{i=1}^{n}y_{i}\\ &=(\sqrt{n}+n)(1-\sum\limits_{i=1}^{n}y_{i})\\ &\geq 0,\\ \end{array}$
which implies that $x=(x_{1},\dots,x_{n})\in\bar{\Delta}$. By computations, we have that for $(y_{1},\dots,y_{n+1})\in\Delta^{n+1}$, the following holds
$\widehat{F}(y_{1},\dots,y_{n+1})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}^{(d)}B_{\alpha}^{(d)}(x)=F(x)\succeq F_{\min}\cdot I_{\ell}.$
Hence, $\widehat{F}\succeq F_{\min}\cdot I_{\ell}$ on $\Delta^{n+1}$. It follows from [48, Theorem 3] that for $k>d(d-1)\frac{\|\widehat{F}\|_{2}}{2F_{\min}}-d$, all coefficient matrices of $(y_{1}+\cdots+y_{n+1})^{k}\widehat{F}(y_{1},\dots,y_{n+1})$ in the monomial basis are positive definite, i.e., there exist $P_{\alpha}\succ 0$ such that
@@
into the above identity, we obtain
@@
Since $\|\widehat{F}\|_{2}=\|F\|_{B}$, the conclusion follows.
@@
Then, we have that $\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}$. Furthermore, there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that $\|P^{(\xi_{0})}\|_{B,d_{P}}=\|P\|_{B}$.
Proof 3.4.
Suppose that $P$ can be expressed in the Bernstein basis $\{B_{\alpha}^{(d_{P})}(x):\alpha\in\mathbb{N}^{n}_{d_{P}}\}$ as
@@
$\begin{array}[]{ll}\|P^{(\xi)}\|_{B,d_{P}}=\|\xi^{T}P(x)\xi\|_{B,d_{P}}&=\|\sum\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\xi^{T}P_{\alpha}^{(d_{P})}\xi B_{\alpha}^{(d_{P})}(x)\|_{B,d_{P}}\\ &=\max\{|\xi^{T}P_{\alpha}^{(d_{P})}\xi|:~{}\alpha\in\mathbb{N}^{n}_{d_{P}}\}\\ &\leq\max\{\|P_{\alpha}^{(d_{P})}\|_{2}:~{}\alpha\in\mathbb{N}^{n}_{d_{P}}\}=\|P\|_{B}.\end{array}$
Thus, we have $\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}.$ Since there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that
$\|P\|_{B}=\max\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|:~{}\alpha\in\mathbb{N}^{n}_{d_{P}}\},$
@@
$\|P^{(\xi_{0})}\|_{B,d_{P}}=\max\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|:~{}\alpha\in\mathbb{N}^{n}_{d_{P}}\}=\|P\|_{B}.$
@@
(3.4) $\mathcal{G}(x)=-\min\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{g_{1}(x)}{\left\|g_{1}\right\|_{B}}},\ldots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{g_{m}(x)}{\left\|g_{m}\right\|_{B}}},0\right).$
@@
If $\mathcal{G}(x)=0$, then we have $x\in K$ and $d(x,K)=0$. The distance function $d(x,K)$ is a continuous semialgebraic function (see [5, Proposition 2.2.8], [14, Proposition 1.8]). By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Łojasiewicz inequality holds:
@@
In the following, we prove a polynomial bound on the degrees of terms appearing in matrix Putinar’s Positivstellensätz when $K$ is given by finitely many scalar polynomial inequalities, which improves the exponential bound established in [15]. Recall that the quadratic module of scalar polynomials generated by $G(x)$ in $\mathbb{R}[x]$ is denoted by $\mbox{QM}[G]$ (see Section 2.2).
@@
Let $G(x)$ be a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$ and let $K$ be as in (3.1). Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{\ell}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
@@
In the above, $\kappa$, $\eta$ are respectively the Łojasiewicz constant and exponent in (3.5) and $C>0$ is a constant independent of $F$, $g_{1},\dots,g_{m}$. Furthermore, the Łojasiewicz exponent can be estimated as $\eta\leq(d_{G}+1)(3d_{G})^{n+m-2}$.
Remark. At a feasible point $x\in K$, the Mangasarian–Fromovitz constraint qualification (MFCQ) holds if there exists a direction $v\in\mathbb{R}^{n}$ such that $\nabla g_{j}(x)^{T}v>0$ for all active inequality constraints $g_{j}(x)=0$. The linear independence constraint qualification (LICQ) holds at $x$ if the set of gradient vectors $\nabla g_{j}(x)$ of all active constraints is linearly independent. It was shown in [45] that if the MFCQ holds at every $x\in K$, then the Łojasiewicz exponent, as in (3.5), is equal to 1. Since the MFCQ is weaker than the LICQ, we know that if the LICQ holds at every $x\in K$, the Łojasiewicz exponent $\eta=1$ (see also [2, 3]) [^1] The author would like to thank the referee for pointing out this fact.. Then, under the assumptions of Theorem 3.5, we have that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F}\in\mbox{QM}[G]_{k}^{\ell}$ if
Proof 3.6.
Without loss of generality, we assume that $\left\|g_{i}\right\|_{B}=1$ for $i=1,\dots,m$. Otherwise, we can scale $g_{i}$ by $1/\|g_{i}\|_{B}$. We remark that scaling $g_{i}$ does not change $\kappa$, since we normalize $g_{i}$ in $\mathcal{G}$ and $(g_{i}(x)/\|g_{i}\|_{B})/(\|g_{i}(x)/\|g_{i}\|_{B}\|_{B})=g_{i}(x)/\|g_{i}\|_{B}$. Thus, the bound remains unchanged. Let
@@
For the above $\delta$, $v$, it follows from [3, Proposition 3.2] that there exists $h_{i}\in\mathbb{R}[x]$ such that for $x\in\bar{\Delta}$, we have that
@@
In the following, we show that $P\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$. Equivalently, for a fixed $x\in\bar{\Delta}$, we prove that for each $\xi\in\mathbb{R}^{\ell}$ with $\|\xi\|_{2}=1$, the following holds: